V=?
F=40kg
M=2kg
t=3s
1
Expert's answer
2018-08-31T15:27:09-0400
Let’s first find the acceleration of the object from the Newton’s Second Law of Motion:
F=Ma,
a=F/M=(40 N)/(2 kg)=20 m/s^2 .
Finally, we can find the final velocity of the object from the kinematic equation:
v=v_0+at=0 m/s+20 m/s^2 ∙3 s=60 m/s.